# Teola: Towards End-to-End Optimization of LLM-based Applications

### v1

@@ 9. Related Work
Frameworks for LLM-based applications. Apart from frameworks like (llamaindex, ; langchain, ; promptflow, ), several studies (kim2023llmcompiler, ; khattab2023dspy, ; zheng2023sglang, ; lin2024parrot, ) focus on optimizing complex LLM tasks involving multiple LLM calls. They explore opportunities such as parallelism and sharing by developing specific programming interfaces or compilers. Moreover, several AI agent frameworks (langgraph, ; wu2023autogen, ; shen2023hugginggpt, ; hong2023metagpt, ; khattab2023dspy, ; hong2024data, ; lin2024parrot, ) enable LLMs to autonomously control workflows, make decisions, select tools, and interact with other LLMs, reducing the need for human intervention but introducing specific challenges. Teola is more similar to (llamaindex, ; langchain, ; promptflow, ), maintaining an app of various components with human-defined flow using a primitive-level graph while focusing on end-to-end execution efficiency. Parrot (lin2024parrot, ) also captures the application-level affinity of multiple LLM requests using prompt structure to facilitate joint scheduling. Orthogonally, Teola focuses on full execution graph optimizations for applications involing both LLM and non-LLM parts.
ML analytics systems. Existing ML analytics systems, such as VideoStorm (zhang2017live, ), Jellybean (wu2022servingmlworkflow, ), Llama (romero2021llama, ), and Vulcan (zhang2024vulcan, ), focus on optimizing video analytics pipelines by configuring and placing components across heterogeneous resources. While they are similar to LLM-based workflows, the latter involves more complex request patterns and flexible configurations. Besides, video systems often maintain uniform pipelines for all queries and overlook coordination between frontend orchestration and backend scheduling. Teola addresses these limitations by constructing finer-grained dataflow graph for orchestration, enhancing execution and scheduling efficiency, and leveraging LLM-specific attributes.

### v2

@@ 9. Related Work
Frameworks for LLM applications. Apart from frameworks mentioned before (llamaindex, ; langchain, ; promptflow, ), several studies (kim2023llmcompiler, ; khattab2023dspy, ; zheng2023sglang, ; lin2024parrot, ) focus on optimizing complex LLM tasks involving multiple LLM calls. They explore opportunities such as parallelism and KV cache sharing by developing specific programming interfaces or compilers. AI agent frameworks (langgraph, ; shen2023hugginggpt, ; hong2023metagpt, ; hong2024data, ) enable LLMs to make decisions, select tools, and interact with other LLMs, reducing the need for human intervention but introducing specific challenges. Teola is more related to (llamaindex, ; langchain, ; promptflow, ), working with an application of various components with a pre-defined workflow while focusing on end-to-end efficiency. Parrot (lin2024parrot, ) also captures the application-level affinity of multiple LLM requests and facilitates joint scheduling. SGlang (zheng2023sglang, ) explores the relationship between LLM requests and attempts to reuse the history KV cache efficiently. They share a similar spirit with us in exploiting opportunities beyond a single inference request. The difference is also clear: Teola focuses on end-to-end optimization at the application level that involve both LLM and non-LLM components, which were not considered before.
Data analytics systems. Data analytics systems (isard2007dryad, ; zaharia2016apachespark, ; gog2015musketeer, ; bauer2012legion, ) employ various graph based dependency analysis to enhance distributed task execution of big data analytics workloads. Teola draws upon these techniques to perform its graph optimization. However, these systems are designed for general-purpose data processing. Teola introduces a new set of abstractions specifically designed for LLM-based workflows, which allows it to explore a much larger optimization space beyond graph optimization or those mentioned in this work.